# README: Bridging Medical Jargon and Lay Understanding for Patient Education through Data-Centric NLP

### v1

@@ Abstract
The advancement in healthcare has shifted focus toward patient-centric approaches, particularly in self-care and patient education, facilitated by access to Electronic Health Records (EHR). However, medical jargon in EHRs poses significant challenges in patient comprehension. To address this, we introduce a new task of automatically generating lay definitions, aiming to simplify complex medical terms into patient-friendly lay language. We first created the README dataset, an extensive collection of over 20,000 unique medical terms and 300,000 mentions, each offering context-aware lay definitions manually annotated by domain experts. We have also engineered a data-centric Human-AI pipeline that synergizes data filtering, augmentation, and selection to improve data quality. We then used README as the training data for models and leveraged a Retrieval-Augmented Generation (RAG) method to reduce hallucinations and improve the quality of model outputs. Our extensive automatic and human evaluations demonstrate that open-source mobile-friendly models, when fine-tuned with high-quality data, are capable of matching or even surpassing the performance of state-of-the-art closed-source large language models like ChatGPT. This research represents a significant stride in closing the knowledge gap in patient education and advancing patient-centric healthcare solutions  [^1] Our code and part of data will be released at https://github.com/seasonyao/NoteAid-README.
@@ 1 Introduction
The advancement in healthcare transcends medical breakthroughs, encompassing an increased emphasis on patient involvement in self-care Boling (2009); Spooner et al. (2019); Baldry et al. (1986); Schillinger et al. (2009). The criticality of effective patient education is underscored in an era marked by the emergence of new treatments that demand patient comprehension and cooperation. Access to Electronic Health Records (EHR) has become an integral part of this educational paradigm, with initiatives like PCASSO Masys et al. (2002) and NoteAid Chen et al. (2018); Kwon et al. (2022) leading the way in empowering patients and enhancing healthcare outcomes. These efforts demonstrate a shift towards patient-centric healthcare, where informed patients play an active role in their treatment processes.
@@
Despite advancements in electronic medical record management, one significant barrier persists in the form of medical jargon in EHRs, impeding patient understanding and self-care Kwon et al. (2022). As shown in Figure 1, tools like NoteAid, which employs Natural Language Processing (NLP) to demystify complex medical terms, have been instrumental in bridging the communication gap between healthcare professionals and patients Lalor et al. (2018, 2021, 2023). However, existing resources like the Consumer Health Vocabulary (CHV) Zeng and Tse (2006); He et al. (2017); Ibrahim et al. (2020) are limited in scale, posing a challenge to NoteAid. For instance, only a fraction (about 4%) of the medical terms in NoteAid have been annotated with lay definitions, highlighting the need for a more scalable solution to address this knowledge gap effectively.
@@
Addressing this issue requires shifting our focus to online health education materials such as Unified Medical Language System (UMLS)  Bodenreider (2004), MedlinePlus Patrias and Wendling (2007), Wikipedia, and Google. However, as indicated in Figure 1, these resources often present information that is too difficult for the average patient to understand. For example, these resources’ average readability measured by the Flesch Kincaid Grade Level was post-secondary or higher education, while the average readability of a US adult was 7-8th grade level Doak et al. (1996, 1998); Eltorai et al. (2014). To bridge this gap, we have engaged a team of medical experts to meticulously curate lay definitions for jargon terms found in NoteAid, targeting a comprehension level suitable for individuals with a 7th to 8th-grade education. Each term within the NoteAid dataset has been redefined across various contexts, ensuring their applicability in diverse clinical scenarios. This effort led to the creation of the REsource of lAy Definitions for MEdical jargon (README) dataset, an expansive resource containing lay definitions for over 20,000 medical terms. In total, the README dataset comprises an impressive 300,000 data points, each consisting of a clinical note context, a medical jargon term, and its corresponding lay definition, thereby significantly enhancing the accessibility and comprehensibility of medical information for patients.
Figure 1: The NoteAid pipeline comprises an NLP component and a resource of medical jargon and the corresponding lay definitions. We list definitions for the term "EGD" across various online sources registered Flesch-Kincaid Grade Level (FKGL) above 13, suggesting they are comprehensible mainly to individuals with post-secondary education.
Figure 2: Our Human-AI-in-the-loop Data-Centric NLP pipeline, comprising the Examiner-Augmenter-Examiner (EAE) framework and different data selection methods. EAE shows how humans (physicians) and AI (LLM, e.g. ChatGPT) cooperate to make a high-quality README dataset. We collect general definitions for every jargon term from external knowledge resources such as UMLS. “R” is “README”. “v1” is “version-1”. “instruction” and “demo” (examples for in-context learning) are combined into the prompt for LLM. In the pipeline, the human duties at different stages are annotator (labeling the initial dataset) and instructor (providing suitable prompts to guide AI at every stage). The AI duties at different stages are examiner (filter high-quality data) and augmenter (improve the quality of low-quality data). We then deploy 4 different data selection strategies to combine high-quality expert-annotated data R-v2 and high-quality AI-synthetic data R-v5 and train the in-house system.
@@
Yet, the critical aspect of generating lay definitions remains largely unexplored. As patients gain more access to their EHRs, the demand for lay definition resources is escalating, inevitably destined to surpass the capacity of current expert-annotated resources, regardless of efforts to expand them. In addition, the dynamic nature of "jargon" based on individual and context makes pre-annotated expert resources less adaptable to real-life scenarios. The model-driven automatic generation of lay definitions from medical jargon emerges as a viable solution. Recent research highlighted ChatGPT’s potential in its integration with the field of medicine Brown et al. (2020); OpenAI (2023); Yang et al. (2023), including generating human-readable definitions for biomedical terms Remy and Demeester (2023). Nonetheless, our evaluations of open-source models (refer to Figure 4) indicate notable differences in their performance compared to ChatGPT, particularly in terms of medical knowledge Sung et al. (2021); Yao et al. (2022, 2023).
@@
To bridge this gap, we aim to train an in-house system using open-source models for automatic lay definition generation to provide reliable lay definitions for jargon in patient education tools like NoteAid. Inspired by research on Retrieval-Augmented Generation (RAG)  Lewis et al. (2020); Asai et al. (2023), we aim to overcome the limitations of open-source base models in medical knowledge. We are positioning automatic lay definition generation as a form of text simplification, where language models are prompted to generate context-aware, jargon-specific, and layperson-friendly definitions based on standard definitions retrieved from external knowledge resources. Specifically, in this work, we use the UMLS to retrieve standard definitions of jargon terms and construct a dataset upon the README that includes context, jargon terms, standard definitions, and lay definitions.
@@
Subsequently, we designed the data-centric pipeline Examiner-Augmenter-Examiner (EAE) to enhance the data quality within the README dataset, as seen in Figure 2. Specifically, following a human-in-the-loop paradigm Monarch (2021), we employed human experts to guide AI in both examiner and augmenter stages, where the former selects high-quality training data (originating from either manual annotation or AI synthesis), and the latter generates potentially high-quality synthesized data to increase data points. Finally, we employed the AI-synthesized dataset to augment the expert-annotated dataset, aiming to explore the effectiveness of AI-generated data in training, especially in scenarios with limited expert data. We implemented a range of heuristic data selection strategies to integrate AI synthetic data, allowing us to incorporate suitable data points into our training process.
@@
  • [leftmargin=.2in,topsep=0.1pt]
  • Introduced a new task of automatically generating lay definitions for medical jargon. We created a substantial expert-annotated dataset containing 300,000 data points, serving as a detailed dictionary and a benchmark for this novel task. We leveraged a Retrieval-Augmented Generation (RAG) method, which aims to reduce hallucinations and improve the quality of model outputs. Our work aims to enhance patient education by improving patients’ comprehension of medical documentation.
@@
2 Related Work
Data-Centric AI In the Data-centric AI framework, the README meticulously navigates through the collection, labeling, preparation, reduction, and augmentation phases Zha et al. (2023); Ng et al. (2021). README begins with data collection and labeling (4.1), harnessing domain expertise to curate a dataset that is both expansive and representative Whang et al. (2023). During the preparation phase (4.2), raw data undergoes rigorous cleaning and transformation, readying it for effective model training Krishnan and Wu (2019). README also adeptly applies data augmentation techniques (4.2) to enrich the dataset with verified quality AI-synthetic data Chlap et al. (2021). Finally, through data reduction (4.3), README selects more suitable instances from AI-synthetic data for data integration with expert-annotated data Whang et al. (2023). Collectively, these stages underscore README’s comprehensive application of Data-centric AI in the healthcare domain.
Patient Education Patient education plays a crucial role in the success of therapy (Bastable, 2016; Golper, 2001). Elevating the level of patient education has consistently been an important component within the healthcare system (McCarthy et al., 2013). By providing clear and easily understandable medical information, patient education not only enhances patients’ awareness of their own health conditions but also encourages them to actively participate in medical decision-making and self-management(Gruman et al., 2010; Coulter, 2012). NoteAid (Chen et al., 2018) transforms complex medical terms in EHR notes into understandable language, integrating advanced features like MedJEx (Kwon et al., 2022) to enhance patient comprehension. Efforts to demystify clinical language have also been made in the works of  Petrova (2014); Petrova et al. (2015), while Remy and Demeester (2023) successfully employed ChatGPT for generating accurate medical definitions.

### v2

@@ Abstract
The advancement in healthcare has shifted focus toward patient-centric approaches, particularly in self-care and patient education, facilitated by access to Electronic Health Records (EHR). However, medical jargon in EHRs poses significant challenges in patient comprehension. To address this, we introduce a new task of automatically generating lay definitions, aiming to simplify complex medical terms into patient-friendly lay language. We first created the README dataset, an extensive collection of over 50,000 unique (medical term, lay definition) pairs and 300,000 mentions, each offering context-aware lay definitions manually annotated by domain experts. We have also engineered a data-centric Human-AI pipeline that synergizes data filtering, augmentation, and selection to improve data quality. We then used README as the training data for models and leveraged a Retrieval-Augmented Generation method to reduce hallucinations and improve the quality of model outputs. Our extensive automatic and human evaluations demonstrate that open-source mobile-friendly models, when fine-tuned with high-quality data, are capable of matching or even surpassing the performance of state-of-the-art closed-source large language models like ChatGPT. This research represents a significant stride in closing the knowledge gap in patient education and advancing patient-centric healthcare solutions   [^1] Our code and data will be released at https://github.com/seasonyao/NoteAid-README.
@@ 1 Introduction
Figure 1: A visualization of the NoteAid pipeline, where NLP tools first identify jargon that may be challenging for patients to understand. The lay definitions corresponding to these jargon terms are then retrieved from relevant dictionaries and presented to the patients, enhancing their comprehension and engagement with their health information.
@@
The advancement in healthcare transcends medical breakthroughs, encompassing an increased emphasis on patient involvement in self-care Boling (2009); Spooner et al. (2019); Baldry et al. (1986); Schillinger et al. (2009). The criticality of effective patient education is underscored in an era marked by the emergence of new treatments that demand patient comprehension and cooperation. Access to Electronic Health Records (EHR) has become a foundational element of modern educational paradigms Gabay (2017); Lye et al. (2018). Despite advancements in electronic medical record management Walker et al. (2019), one significant barrier persists in the form of medical jargon   [^2] A “medical jargon term” is the specialized language healthcare professionals use that can be complex for non-medical individuals. A “lay definition” translates this jargon into accessible language for the general public, aiming to bridge the understanding gap and enhance the communication of health-related information. in EHRs, impeding patient understanding and self-care Kujala et al. (2022); Choudhry et al. (2016); Khasawneh et al. (2022). As shown in Figure 1, tools like NoteAid Chen et al. (2018); Kwon et al. (2022), which employs Natural Language Processing (NLP) to demystify complex medical terms, have been instrumental in bridging the communication gap between healthcare professionals and patients Lalor et al. (2018, 2021, 2023). However, lay language dictionary resources like the Consumer Health Vocabulary (CHV) Zeng and Tse (2006); He et al. (2017); Ibrahim et al. (2020) are limited in scale, posing a challenge to NoteAid. For instance, only a fraction (about 4%) of the medical terms in NoteAid have been annotated with lay definitions, highlighting the need for a more scalable solution to address this knowledge gap effectively.
@@
Addressing this issue requires shifting our focus to online health education materials such as Unified Medical Language System (UMLS)  Bodenreider (2004), MedlinePlus Patrias and Wendling (2007), Wikipedia, and Google. However, as indicated in Appendix Table 4, these resources often present too difficult information for the average patient to understand. For example, these resources’ average readability measured by the Flesch Kincaid Grade Level Solnyshkina et al. (2017) was post-secondary or higher education, while the average readability of a US adult was 7-8th grade level Doak et al. (1996, 1998); Eltorai et al. (2014). To bridge this gap, we have engaged medical experts to meticulously curate lay definitions for jargon terms found in NoteAid-MedJEx Kwon et al. (2022), targeting a comprehension level suitable for individuals with a 7th to 8th-grade education. Each jargon term has been redefined across various contexts, ensuring their applicability in diverse clinical scenarios. This effort led to our creation of the REsource of lAy Definitions for MEdical jargon (README) dataset, an expansive resource containing over 51,623 (medical jargon term, lay definition) pairs. The README dataset comprises an impressive 308,242 data points, each consisting of a clinical note context, a medical jargon term, and its corresponding lay definition. Thus, the dataset significantly enhances the accessibility and comprehensibility of medical information for patients.
@@
Yet, the critical aspect of generating lay definitions remains largely unexplored. As patients gain more access to their EHRs, the demand for lay definition resources is escalating. Despite our efforts to expand them, they are inevitably destined to surpass the capacity of current expert-annotated resources. Moreover, the dynamic nature of "jargon" based on individual and context makes pre-annotated expert resources less adaptable to real-life scenarios. The model-driven automatic generation of lay definitions from medical jargon emerges as a viable solution. Recent research highlighted ChatGPT’s potential in its integration with the field of medicine Brown et al. (2020); OpenAI (2023); Yang et al. (2023), including generating human-readable definitions for biomedical terms Remy and Demeester (2023). Nonetheless, our evaluation of open-source models (refer to Figure 3) shows a significant performance degradation compared to ChatGPT. Using open-source large language models like Llama2 Touvron et al. (2023) and small language models such as GPT-2 Radford et al. (2019) is crucial because proprietary LLMs accessed via third-party APIs may not always be feasible, especially in fields like healthcare with strict privacy requirements and economic constraints. Open-source models offer the necessary privacy, while smaller models provide economic and infrastructural benefits, addressing distinct concerns about effectively deploying NLP tools in healthcare scenarios.
@@
To bridge this gap, we aim to train an in-house system using open-source models for automatic lay definition generation to provide reliable lay definitions for jargon in patient education tools like NoteAid. Inspired by research on Retrieval-Augmented Generation (RAG) in general and medical domains Lewis et al. (2020); Asai et al. (2023); Xiong et al. (2024); Wang et al. (2024); Guo et al. (2024), we designed to use external resources to overcome the limitations of these open-source models in medical knowledge Sung et al. (2021); Yao et al. (2022, 2023); Chen et al. (2023). We are positioning automatic lay definition generation as a form of text simplification, where language models are prompted to generate context-aware, jargon-specific, and layperson-friendly definitions based on general definitions retrieved from external knowledge resources. Specifically, in this work, we use the UMLS to retrieve general definitions of jargon terms and construct a dataset upon the README that includes context, jargon terms, general definitions, and lay definitions.
@@
To improve the initial README dataset’s data quality for model training, we developed a data-focused process called Examiner-Augmenter-Examiner (EAE), as illustrated in Figure 2. Drawing inspiration from the human-in-the-loop concept Monarch (2021), we employed human experts to guide AI in both the examiner and augmenter stages. The examiner filters high-quality training data, which may come from expert annotations before the augmenter stage or from AI-generated content after. The augmenter generates potentially high-quality synthesized data to increase data points. In the end of EAE, Human annotators review filtered data to ensure its quality. After obtaining high-quality expert-annotated and AI-synthesized datasets, we employed the AI-synthesized dataset to augment the expert-annotated dataset, aiming to explore the effectiveness of AI-synthesized data in training. We implemented a range of heuristic data selection strategies to integrate AI synthetic data, allowing us to incorporate suitable data points into our training process.
@@
  • Introduced a new task of automatically generating lay definitions for medical jargon. We created a substantial expert-annotated dataset of 308,242 data points that can be used directly as a detailed lay-language dictionary for patient education tools such as NoteAid and as training data for this new task. We leveraged a Retrieval-Augmented Generation (RAG) method to reduce hallucinations and improve the quality of model outputs. Our work aims to enhance patient education by improving patients’ comprehension of medical documentation.
@@ 2 Problem Statement
Figure 2: Our Data-Centric NLP pipeline, comprising the Examiner-Augmenter-Examiner (EAE) framework and different data selection methods. EAE shows how humans (physicians) and AI (LLM, e.g. ChatGPT) cooperate to make a high-quality README dataset. We collect general definitions for every jargon term from external knowledge resources such as UMLS. “R” is “README”. “exp” is “expert annotation version”, “syn” is “AI synthetic version”. “instruction” and “demo” (examples for in-context learning) are combined into the prompt for LLM. In the pipeline, the human duties at different stages are annotator (labeling the initial dataset) and instructor (providing suitable prompts to guide AI at every stage). The AI duties at different stages are examiner (filter high-quality data) and augmenter (improve the quality of low-quality data). Appendix Table 5 describes the number of different versions of the README dataset in each step. After we get two high-quality datasets, R-exp_good and R-syn_good, from the EAE pipeline, we then deploy 4 different data selection strategies to combine high-quality expert-annotated data R-exp_good and high-quality AI-synthetic data R-syn_bad for in-house system training.